# Track Anything Rapter(TAR)

### v1

@@ Abstract
Object tracking is a fundamental task in computer vision with broad practical applications across various domains, including traffic monitoring, robotics, and autonomous vehicle tracking. In this project, we aim to develop a sophisticated aerial vehicle system known as Track Anything Raptor (TAR), designed to detect, segment, and track objects of interest based on user-provided multimodal queries, such as text, images, and clicks. TAR utilizes cutting-edge pre-trained models like DINO, CLIP, and SAM to estimate the relative pose of the queried object. The tracking problem is approached as a Visual Servoing task, enabling the UAV to consistently focus on the object through advanced motion planning and control algorithms. We showcase how the integration of these foundational models with a custom high-level control algorithm results in a highly stable and precise tracking system deployed on a custom-built PX4 Autopilot-enabled Voxl2 M500 drone. To validate the tracking algorithm’s performance, we compare it against Vicon-based ground truth. Additionally, we evaluate the reliability of the foundational models in aiding tracking in scenarios involving occlusions. Finally, we test and validate the model’s ability to work seamlessly with multiple modalities, such as click, bounding box, and image templates.
@@ 1 Introduction
Figure 2: Rapter: The custom built drone used for the experiments
@@ 1.1 Contribution
    ROS2-Based Implementation: We implemented an equivalent ROS2-based system utilizing MAVROS, thereby bypassing the need for MAVSDK prposed in the base paper for drone control, facilitating enhanced integration and operational flexibility.
@@
    DTW based Trajectory evaluation:Finally, we discussed the utilization of standard metrics such as Discrete Time Warping (DTW) to evaluate the trajectory traced by the drone during object tracking. This approach provides a systematic method for assessing the quality and effectiveness of the proposed tracking algorithm. By employing DTW, we can quantitatively measure the similarity between the drone’s trajectory and the expected path, providing valuable insights into the performance and accuracy of the tracking system.
@@ 2.1 Detection
In our pipeline, we utilize SAM and DINO models for object detection and segmentation. The input comprises an RGB frame $F$ from a video stream and a query $q$, which can be in the form of an image template, click or a bounding box on the input image.
@@ 2.2 Re-detecting a lost object
Automatic re-detection via cross trajectory stored ViT features: To enable a robust and accurate autonomous re-detection of the tracked (lost) object, we provide a feature-descriptor storing mechanism for the tracked object in different stages of the tracking process, these stored features will be used to find the object once lost. Specifically, we suggest the following. Let $\tau>0$ be an integer. During the tracking, at each iteration $i$ such that $i\mod\tau=0$, define $M_{obj}^{i}$ to be the mask denoting the current tracked object in the frame, we first apply $Desc$ on the current frame $F_{i}$ to obtain $D_{i}:=Desc(F_{i})\in\mathbb{R}^{h\times w\times d}$, then we compute the mean descriptor of the current tracked object as:
@@ 3.2 Simulation
The implemented tracking algorithm was evaluated in a ROS2 Gazebo environment, where a PX4-based simulator drone was tasked with tracking an Apriltag in motion. The tracking was performed on-the-fly, based on a bounding box query provided around the Apriltag model, without any prior training. The results can be seen in 6.
@@ 3.3.2 Process Flow
The processing of the incoming video feed and the generation of appropriate control signals are achieved using two ROS2 nodes as depicted in the figure 4. The first node, named the Vision node, is responsible for processing the incoming visual feed to generate the appropriate control signals. The second node, named the Controller node, is responsible for utilizing the low-level controller to execute motions based on the obtained control signals.
The Vision node is designed to control the UAVs drone’s movement based on the position of a tracked object within a video frame. The core principle involves calculating a direction vector,
@@
which determine the drone’s velocity and heading adjustments. To enhance robustness, a filtering process utilizing the coefficients $\mathbf{b}$ and $\mathbf{a}$ is applied to smooth the control signals, mitigating the effects of noise and abrupt positional changes. This results in a filtered direction vector,
@@
Figure 6: Oneshot Apriltag tracking based on the realtime bounding box query.
@@ 4.2 Multimodal Input Queries
These offsets were significantly reduced when DINO was introduced to detect and track the features. In Figure 15, it is evident that the Raptor’s trajectory is more stable compared to previous figures. This is corroborated by the calculated DTW for this run, which shows a mean DTW of 0.67 meters, indicating the drone had an offset of only 0.67 meters.
To further verify DINO’s capabilities, we switched the target from an AprilTag back to a toy turtle and provided an image template of the turtle. DINO effortlessly detected the features and provided similar results to those obtained with the AprilTag. As shown in Figure 17, the mean DTW is 0.62 meters, a remarkably low offset considering the specifications of the Raptor’s tracking camera.
@@ 4.4 Tracking against obstruction
To test the robustness of the Raptor’s multi-modality, we conducted several test runs introducing occlusion to the target, tracking it while gradually occluding it from one end of the setup and re-emerging it from the other, ensuring the entire setup was within the Raptor’s field of view (FOV). While testing on a wide dataset of frames, the Raptor could easily re-detect the target using DINO but failed to re-detect the targets on its own when they re-emerged. To test if it could detect any features, we assisted the Raptor by switching to ”position mode” mid-run and flew it slightly closer to the target until it could re-detect it, which the Raptor successfully did. After numerous runs and analyses, we concluded that the re-detection algorithm itself is robust, but limited hardware and external conditions, which will be discussed later, inhibited the drone from re-detecting the target on its own. In Figures 19 and 21, a peculiar pattern in the Raptor’s trajectory indicates where it lost visuals on the target.

### v2

@@ Abstract
Object tracking is a fundamental task in computer vision with broad practical applications across various domains, including traffic monitoring, robotics, and autonomous vehicle tracking. In this project, we aim to develop a sophisticated aerial vehicle system known as Track Anything Rapter (TAR), designed to detect, segment, and track objects of interest based on user-provided multimodal queries, such as text, images, and clicks. TAR utilizes cutting-edge pre-trained models like DINO, CLIP, and SAM to estimate the relative pose of the queried object. The tracking problem is approached as a Visual Servoing task, enabling the UAV to consistently focus on the object through advanced motion planning and control algorithms. We showcase how the integration of these foundational models with a custom high-level control algorithm results in a highly stable and precise tracking system deployed on a custom-built PX4 Autopilot-enabled Voxl2 M500 drone. To validate the tracking algorithm’s performance, we compare it against Vicon-based ground truth. Additionally, we evaluate the reliability of the foundational models in aiding tracking in scenarios involving occlusions. Finally, we test and validate the model’s ability to work seamlessly with multiple modalities, such as click, bounding box, and image templates.
@@ 1 Introduction
Figure 2: Rapter: The custom-built drone used for the experiments
@@ 1.1 Contribution
    ROS2-Based Implementation: We implemented an equivalent ROS2-based system utilizing MAVROS, thereby bypassing the need for MAVSDK proposed in the base paper for drone control, facilitating enhanced integration and operational flexibility.
@@
    DTW-based Trajectory evaluation:Finally, we discussed the utilization of standard metrics such as Discrete Time Warping (DTW) to evaluate the trajectory traced by the drone during object tracking. This approach provides a systematic method for assessing the quality and effectiveness of the proposed tracking algorithm. By employing DTW, we can quantitatively measure the similarity between the drone’s trajectory and the expected path, providing valuable insights into the performance and accuracy of the tracking system.
@@ 2.1 Detection
In our pipeline, we utilize SAM and DINO models for object detection and segmentation. The input comprises an RGB frame $F$ from a video stream and a query $q$, which can be in the form of an image template, click, or a bounding box on the input image.
@@ 2.2 Re-detecting a lost object
Automatic re-detection via cross-trajectory stored ViT features: To enable a robust and accurate autonomous re-detection of the tracked (lost) object, we provide a feature-descriptor storing mechanism for the tracked object in different stages of the tracking process, these stored features will be used to find the object once lost. Specifically, we suggest the following. Let $\tau>0$ be an integer. During the tracking, at each iteration $i$ such that $i\mod\tau=0$, define $M_{obj}^{i}$ to be the mask denoting the current tracked object in the frame, we first apply $Desc$ on the current frame $F_{i}$ to obtain $D_{i}:=Desc(F_{i})\in\mathbb{R}^{h\times w\times d}$, then we compute the mean descriptor of the current tracked object as:
@@ 3.2 Simulation
The implemented tracking algorithm was evaluated in a ROS2 Gazebo environment, where a PX4-based simulator drone was tasked with tracking an Apriltag in motion. The tracking was performed on the fly, based on a bounding box query provided around the Apriltag model, without any prior training. The results can be seen in 6.
@@ 3.3.2 Process Flow
The processing of the incoming video feed and the generation of appropriate control signals are achieved using two ROS2 nodes as depicted in figure 4. The first node, named the Vision node, is responsible for processing the incoming visual feed to generate the appropriate control signals. The second node, named the Controller node, is responsible for utilizing the low-level controller to execute motions based on the obtained control signals.
The Vision node is designed to control the UAV drone’s movement based on the position of a tracked object within a video frame. The core principle involves calculating a direction vector,
@@
which determines the drone’s velocity and heading adjustments. To enhance robustness, a filtering process utilizing the coefficients $\mathbf{b}$ and $\mathbf{a}$ is applied to smooth the control signals, mitigating the effects of noise and abrupt positional changes. This results in a filtered direction vector,
@@
Figure 6: Oneshot Apriltag tracking based on the real-time bounding box query.
@@ 4.2 Multimodal Input Queries
These offsets were significantly reduced when DINO was introduced to detect and track the features. In Figure 15, it is evident that the Rapter’s trajectory is more stable compared to previous figures. This is corroborated by the calculated DTW for this run, which shows a mean DTW of 0.67 meters, indicating the drone had an offset of only 0.67 meters.
To further verify DINO’s capabilities, we switched the target from an AprilTag back to a toy turtle and provided an image template of the turtle. DINO effortlessly detected the features and provided similar results to those obtained with the AprilTag. As shown in Figure 17, the mean DTW is 0.62 meters, a remarkably low offset considering the specifications of the Rapter’s tracking camera.
@@ 4.4 Tracking against obstruction
To test the robustness of the Rapter’s multi-modality, we conducted several test runs introducing occlusion to the target, tracking it while gradually occluding it from one end of the setup and re-emerging it from the other, ensuring the entire setup was within the Rapter’s field of view (FOV). While testing on a wide dataset of frames, the Rapter could easily re-detect the target using DINO but failed to re-detect the targets on its own when they re-emerged. To test if it could detect any features, we assisted the Rapter by switching to ”position mode” mid-run and flew it slightly closer to the target until it could re-detect it, which the Rapter successfully did. After numerous runs and analyses, we concluded that the re-detection algorithm itself is robust, but limited hardware and external conditions, which will be discussed later, inhibited the drone from re-detecting the target on its own. In Figures 19 and 21, a peculiar pattern in the Rapter’s trajectory indicates where it lost visuals on the target.